Method of Substitution=Problem Solving
6. Samantha works twice as many hours per week as Adriana. Together they work a total of 39 hour one week.
a) State how ou will assign variables.
b) Write an equation to represent the information in the first sentence.
c) Write an equation to represent the information in the second sentence.
d) Use the method of substitution to find the number of hours worked by each person.
Please and thank you very much.
i tried so much but i didnt get it at all.:(

Answer by tvandenberg(45)

a) Just assign some variables. Let's choose S to be the hours for Samantha and A for Adrianna.
b) S = 2 * A
c) S + A = 39
d) Substituting (b) into (c)... 2 * A + A = 39 .... 3 A = 39 ... A = 13
Substituting back into (b) .... S = 2 * 13 = 26

Hope this helps, Troy.